Circular No. 7644

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001cn IN IC 4758 P. Candia, R. C. Smith, and N. Suntzeff, Cerro Tololo Interamerican

Observatory (CTIO), report: “A spectrum taken by D. Norman and K. Olsen (CTIO) on June 14.3 UT with the Cerro Tololo 1.5-m telescope (+ Cassegrain Spectrograph) shows SN 2001cn (cf. IAUC 7643) to be a type-Ia supernova near maximum light. The supernova photospheric expansion velocity with respect to the host galaxy is 11000 km/s, based on the difference between the NASA/IPAC Extragalactic Database recession velocity for IC 4758 of 4647 km/s and the measured Si II absorption trough (rest 635.5 nm).”

XTE J1859+226 A. V. Filippenko and R. Chornock, University of California at Berke-

ley, write: “Ten spectra (range 400–1000 nm resolution 70 km/s) of XTE J1859+226 (R ∼ 23 mag; IAUC 7506), each of duration 2600–3000 s, were obtained with the Keck-II 10-m telescope (+ Echellette Spectrograph and Imager) on May 29 and 30 UT. The spectra are characteristic of lowmass x-ray binaries, with broad Hα emission superposed on a stellar-plus-accretion-disk continuum. Preliminary radial-velocity measurements for the secondary star were obtained by cross-correlating the individual spectra (range 490–625 nm) against the spectra of several bright dwarfs (G5–K0; G5 gave the best results). The ten velocities provide good phase coverage of a sinusoidal variation, with a velocity amplitude of K = 570 ± 27 km/s, an orbital period of P = 9.16 ± 0.08 hr, and a time of maximum radial velocity at 2001 May 30.440 ± 0.012. This implies a mass function f(M) = 7.4 ± 1.1 solar masses, the largest value ever found (cf. IAUC 7542), strongly suggesting that the compact object is a black hole. The derived period is consistent with values found earlier from photometric observations when the source was bright (P = 9.15 ± 0.05 hr; IAUC 7388) and when it was near quiescence (P = 9.12 ± 0.24 hr; ellipsoidal modulations; IAUC 7506). A second, less-plausible orbital period from our spectroscopic observations is 14.1 hr, but the fit implies an unreasonably high systemic radial velocity (150 km/s) and a very large velocity amplitude (K = 830 km/s); we favor the shorter period and smaller amplitude given above. A more detailed analysis is in progress.”

2001 June 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green